The post office charges $1 to mail an iguana overseas. At this price 1,000 iguana are mailed each week. To increase revenue, the post office is considering an increase in the price of overseas iguana delivery to $2. Post office researchers have determined that the own price elasticity of demand for its iguana delivery service is 1/5. If the post office increases its price to $2: The answer states it will increase by 600. That was the whole question. I am not sure how they got this answer.
Answer #1

Ans: The revenue will increase by 600.

Explanation:

Ep = 1/5 or 0.2

P = $1 , Q = 1000

∆P = $2 -$1 = 1

∆Q = ?

Price elasticity of demand :

Ep = ∆Q/∆P *( P / Q)

0.2 = ∆Q/ 1  *( 1 / 1000)

∆Q = 0.2 *1 *1000 = 200

Since price has increased , quantity demand will fall by 200.

So new quantity = 1000 - 200 = 800

Total revenue before increase in price = Price * Quantity

=$1 *1000 = $1000

Total revenue after  increase in price = Price * Quantity

=$2 * 800 = $1600

Difference in revenue = $1600 - $1000 = $600

So , with the increase in price of post office charges total revenue will increase by 600.